OpenStudy (anonymous):

how do you convert decimal into fraction ? ex 1.5

OpenStudy (unklerhaukus):

well the .5 bit the decimal part of the number is really 5/10 . Decimal means each place behind the dot is 1/10 so you have 1.5 = 1+5/10=1+1/2 =2/2+1/2=3/2

OpenStudy (anonymous):

so what if i had 3.55555

OpenStudy (unklerhaukus):

=3 +5/10 +5/100+5/1000+5/10000+5/100000 =3+55555/100000

OpenStudy (anonymous):

ty